how much pure acid should be mixed with 5 gallons of a 60% acid solution in order to get a 80% acid solution
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of pure acid added (note: that is an 100% solution)
.60*5gal + 1.00*x = .80(5gal + x)
solve for x
.20x = .20*5gal
x = 5 gal of pure acid needs to be added.
checking our answer
.60*5gal + 5gal = 3gal + 5gal = .80*10gal = 8 gal